Circular No. 8259

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003la IN MCG +10-15-89 Further to IAUC 8257, T. Boles reports the discovery of an apparent

supernova (mag 18.0) on unfiltered CCD images taken on Dec. 18.031 and 20.868 UT. SN 2003la is located at α = 10^{h}30^{m}20^{s}.21, δ = +61^{o}15′48′′.5, which is ≈ 4′′.9 west and 1′′.2 south of the center of MCG +10-15-89. SN 2003la is not present on Boles’ images from Feb. 22 and May 6 (limiting mag 19.5) or on Digitized Sky Survey red (1991) or blue (1996) plates.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, write that a spectrum (range 370–750 nm) of SN 2003la, obtained by P. Berlind on Dec. 22.50 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 9233 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 8200 km/s.

SUPERNOVAE 2003ky AND 2003kz Matheson et al. add that a spectrum of SN 2003ky (cf. IAUC 8254),

obtained by M. Calkins on Dec. 18.54 UT, shows it also to be a type-II supernova. Cross-correlation of a spectrum of nearby-host-galaxy light with absorption-line velocity templates indicates a recession velocity of 14100 km/s. Using this value, the expansion velocity derived from the minimum of the Hβ line is 7000 km/s. A spectrum of SN 2003kz (cf. IAUC 8257), obtained by J. Huchra on Dec. 18.54 with the same instrumentation, shows it to be a type-Ia supernova a few weeks past maximum. Cross-correlation of a separate spectrum of the host galaxy nucleus with absorption-line velocity templates indicates a recession velocity of 15100 km/s.

SUPERNOVA 2003jb IN IC 1065 A. V. Filippenko, R. Chornock, and R. J. Foley, University of Califor-

nia, Berkeley, report that inspection of CCD spectra (range 310–1000 nm), obtained on Dec. 22 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2003jb (IAUC 8232) is of type Ia, 1.5–2 months after maximum brightness.

CORRIGENDA On IAUC 8120, ‘Supernovae 2002km–2002ky’, the designations of the

supernovae in the text were given inadvertently as 2003 designations.

2003 December 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green